Using Maxwell’s relations, deduce first and second TdS equations


1
Expert's answer
2021-03-14T19:48:11-0400

Answer

According to thermodynamics

We can express entropy in terms of any two of PVT. Let us first express entropy as a function of V and T

"d S=\left(\frac{\partial S}{\partial V}\right)_{T} d V+\left(\frac{\partial S}{\partial T}\right)_{V} d T."

As

"T d S=T\left(\frac{\partial S}{\partial V}\right)_{T} d V+T\left(\frac{\partial S}{\partial T}\right)_{V} d T."

Now using maxwell equation


"\left(\frac{\partial S}{\partial V}\right)_{T}=\left(\frac{\partial P}{\partial T}\right)_{V}" Using Tds=dU


"T\left(\frac{\partial S}{\partial T}\right)_{V}=\left(\frac{\partial U}{\partial T}\right)_{V}=C_{V}"

This is the first of the TdS equations.


Now let us express entropy as a function of P and T

"d S=\left(\frac{\partial S}{\partial P}\right)_{T} d P+\left(\frac{\partial S}{\partial T}\right)_{P} d T."

"T d S=T\left(\frac{\partial S}{\partial P}\right)_{T} d P+T\left(\frac{\partial S}{\partial T}\right)_{P} d T."

Using maxwell equation

"\left(\frac{\partial S}{\partial P}\right)_{T}=-\left(\frac{\partial V}{\partial T}\right)_{P}"

We know Tds=dU


"T\left(\frac{\partial S}{\partial T}\right)_{P}=\left(\frac{\partial H}{\partial T}\right)_{P}=C_{P}"

So

"T d S=-T\left(\frac{\partial V}{\partial T}\right)_{P} d P+C_{P} d T."

This is the first of the TdS equations.